# Normalized mutual information is a biased measure for classification and community detection

### v1

@@ III.2 Comparison between community detection algorithms
Figure 5 shows results for all six algorithms listed in Section III. Examining the figure we see that in general the best-performing methods are InfoMap, traditional modularity maximization with $\gamma=1$, and the inference method using the degree-corrected stochastic block model. Among the algorithms considered, InfoMap achieves the highest mutual information scores for lower values of the mixing parameter $\mu$ in the LFR model, but fails abruptly as $\mu$ increases, so that beyond a fairly sharp cutoff around $\mu=0.5$ other algorithms do better. InfoMap’s specific failure mode is that it places all nodes in a single community and this behavior can be used as a simple indicator of the failure regime. In this regime one must use another algorithm. Either modularity maximization or inference using the degree-corrected stochastic block model are reasonable options, but modularity has a slight edge, apart from a thin band of intermediate $\mu$ values which, in the interests of simplicity, we choose to ignore. (We discuss some caveats regarding the relationship between the degree-corrected stochastic block model and the LFR benchmark in Appendix A.)
@@ Acknowledgements.
This work was supported in part by the US National Science Foundation under grant DMS–2005899 and by computational resources provided by the Advanced Research Computing initiative at the University of Michigan.
@@ A.2 LFR network generation
The LFR networks we use for benchmarking are generated using the procedure described in LFR08 , which is as follows.
@@
    Fix the number of nodes $n$ and mixing parameter $\mu$. In our studies we use node counts in the range $n\in[200,51200]$. The parameter $\mu$ controls the relative number of edges within and between communities. For small $\mu$ there are many more edges within communities than between them, which makes the communities relatively easy to detect. But as $\mu$ gets larger there are more edges between communities and detection becomes more difficult in a manner reminiscent of the detectability threshold in the standard stochastic block model DKMZ11a ; DKMZ11b . In our studies we use values of $\mu$ in the range $[0.2,0.8]$.
@@
    Draw a degree sequence from a power-law distribution with exponent $\tau_{1}$. Many networks have power-law degree distributions, typically with exponents between 2 and 3 Caldarelli07 , and the LFR model exclusively uses power-law distributions. We use $\tau_{1}=2.5$, with average degree $\langle k\rangle=20$ and maximum degree (which depends on graph size) $k_{\text{max}}=n/10$. Empirically, however, our results do not seem to be very sensitive to these choices.
@@
    Draw a set of community sizes from a power-law distribution with exponent $\tau_{2}$. Many networks are also found to have community sizes that approximately follow a power law, with typical exponents in the range from 1 to 2 Guimera03 ; CNM04 ; PDFV05 ; LFR08 . We use $\tau_{2}=1.5$ with a minimum community size of $s_{\text{min}}=20$ in all cases, while the maximum community size is set to $s_{\text{max}}=\max\left\{n/10,100\right\}$. Again, results were not particularly sensitive to these choices, provided they produce a valid distribution at all.
@@
The parameter values above are similar to those used for instance in YAT16 . As in that study, we find that algorithm performance is dictated primarily by the parameters $n$ and $\mu$, so it is these parameters that are varied our summary figures.
The LFR model is similar to a special case of the degree-corrected stochastic block model (DCSBM) KN11a , and hence one might expect that inference-based community detection methods employing the latter model would perform optimally on LFR networks. Specifically, in the limit of an infinite number of sampled networks and perfect optimization of each community detection method, the final performance measure for any algorithm is given by the expectation value of the similarity $M(g,h[A])$ between the ground truth LFR partition $g$ and the partition $h[A]$ of the LFR network $A$ inferred using the algorithm, where the expectation is taken over the ensemble $P(A,g|\theta)$ of LFR networks and partitions $A,g$ generated using parameters $\theta$ (meaning $\mu$, $\tau_{1}$, $\tau_{2}$, etc.). By using the LFR benchmark with parameters $\theta$ to compare the performance of community detection algorithms, we are therefore implicitly defining the “best” algorithm to be the one whose corresponding function $h[A]$ optimizes $\sum_{A,g}P(A,g|\theta)M(g,h[A])$. If the similarity measure we choose is the “all or nothing” error function $M(g,c)=\delta(g,c)$, then the optimal community detection algorithm is trivially the one with
@@
In other words, the optimal algorithm simply performs maximum a posteriori estimation under the model from which the network was generated. There is no explicit formula for the posterior probability under the LFR model, but to the extent that it is a special case of the DCSBM, we might expect the DCSBM (with appropriate priors) to give optimal results peixoto2021revealing . The LFR model, however, is not precisely a special case of the DCSBM. In particular, the DCSBM normally assumes a uniform distribution over community sizes, where the LRF model assumes a power law. Moreover, we are not using the crude all-or-nothing error function: our entire purpose in this paper is to develop mutual information measures that aggregrate and weigh different modes of error in a sensible fashion. These differences, it appears, are enough to ensure that the DCSBM does not perform the best in our testing.
Regardless, we emphasize that our use of the LFR benchmark in our analysis is simply for consistency with previous studies of network community detection methods LFR08 ; LF09 ; YAT16 . The justification for our proposed similarity measure, on the other hand, is chiefly its theoretical merit over the conventional (symmetric, non-reduced) normalized mutual information, and is independent of the use of the LFR (or any other) benchmark.
Appendix B Upper bound on the
reduced mutual information
In this appendix, we provide evidence for the inequality
$I(c;g)\leq I(g;g),$ (34)
with the exact equality applying if and only if the labelings $c$ and $g$ are identical up to a permutation of the label names. If true, this inequality would imply that the reduced normalized measure $I^{(A)}(c;g)$ was bounded above by 1, as is the unreduced version $I_{0}(c;g)$. We do not at present know of a general proof of (34), but here we present some weaker theoretical results along with strong numerical evidence of its correctness.
First, we note that the only possible violations of (34) would be for small $n$. As $n$ becomes large, the second (reduction) term in Eq. (9) scales as $\log n$, a subleading correction, and hence is negligible for large $n$ compared to the leading mutual information itself which is of order $n$, and the mutual information alone satisfies (34). We have exhaustively checked the following cases with $n$ up to 200 and find (34) to be satisfied in all of them:
$q_{c}=2,\ q_{g}=2$ for $n\leq 200$,
$q_{c}=3,\ q_{g}=2$ for $n\leq 50$,
$q_{c}=4,\ q_{g}=2$ for $n\leq 30$,
$q_{c}=3,\ q_{g}=3$ for $n\leq 30$,
$q_{c}=4,\ q_{g}=3$ for $n\leq 20$.
We also observe zero violations of (9) in any of the (tens of thousands of) randomly generated examples which went into the creation of Fig. 5.
On the theoretical side, (34) implies that $c=g$ is a global maximum of $I(c;g)$, which we are unable to prove, but we can prove the weaker result that it is a local maximum under the mislabeling of any single node, when the reduction term is calculated using Eq. (18). To do this, we first assume, without loss of generality, that a single node with ground truth community label 1 is mislabeled in the candidate partition $c$ as being in community 2. This implies that $n_{1}^{(c)}=n_{1}^{(g)}-1$, $n_{2}^{(c)}=n_{2}^{(g)}+1$, and $n_{r}^{(c)}=n_{r}^{(g)}$ for all $r=3\ldots q_{g}$. The contingency table will then be nearly diagonal, with $(n_{1}^{(g)}-1,n_{2}^{(g)},\ldots)$ on the diagonal and a single nonzero off-diagonal element $n_{12}^{(cg)}=1$. The decrease in the standard mutual information is then given by
$I_{0}(g;g)-I_{0}(g;c)=\log\frac{n!}{\prod_{s}n_{s}^{(g)}!}$
$\quad{}-\log\frac{n!1!(n_{1}^{(g)}-1)!\prod_{s=2}^{q_{g}}n_{s}^{(g)}!}{\left(\prod_{s}n_{s}^{(g)}!\right)\left(\prod_{s=3}^{q_{g}}n_{s}^{(g)}!\right)(n_{1}^{(g)}-1)!(n_{2}^{(g)}+1)!}$
$=\log(n_{2}^{(g)}+1).$ (35)
At the same time, the reduction term $\log\Omega$ also decreases, by an amount that can be estimated from Eq. (18) thus:
$\log$ $\Omega(n^{(g)},n^{(g)})-\log\Omega(n^{(g)},n^{(c)})=\log\prod_{r=1}^{q_{g}}\frac{\binom{n_{r}^{(g)}+\alpha-1}{\alpha-1}}{\binom{n_{r}^{(c)}+\alpha-1}{\alpha-1}}$
$=\log\frac{(n_{1}^{(g)}+\alpha+1)(n_{2}^{(g)}+1)}{n_{1}^{(g)}(n_{2}^{(g)}+\alpha)}.$ (36)
Combining Eqs. (35) and (36), we now find that the change in the RMI incurred from the mislabeling of a single node is
$I(g;g)-I(g;c)=\log\frac{n_{1}^{(g)}(n_{2}^{(g)}+\alpha)}{n_{1}^{(g)}+\alpha-1}>0,$ (37)
where the inequality holds so long as $n_{1}^{(g)},n_{2}^{(g)},\alpha\geq 1$. Thus, under the estimate of $\Omega(n^{(g)},n^{(c)})$ given by Eq. (18), $g=c$ is a local maximum of the reduced mutual information.

### v2

@@ III.2 Comparison between
Figure 6 shows results for all six algorithms listed in Section III. Examining the figure, we see that in general the best-performing methods are InfoMap, traditional modularity maximization with $\gamma=1$, and the inference method using the degree-corrected stochastic block model. Among the algorithms considered, InfoMap achieves the highest mutual information scores for lower values of the mixing parameter $\mu$ in the LFR model, but fails abruptly as $\mu$ increases, so that beyond a fairly sharp cutoff around $\mu=0.5$ other algorithms do better. As noted by previous authors [33], InfoMap’s specific failure mode is that it places all nodes in a single community and this behavior can be used as a simple indicator of the failure regime. In this regime one must use another algorithm. Either modularity maximization or inference using the degree-corrected stochastic block model are reasonable options, but modularity has a slight edge, except in a thin band of intermediate $\mu$ values which, in the interests of simplicity, we choose to ignore. (We discuss some caveats regarding the relationship between the degree-corrected stochastic block model and the LFR benchmark in Appendix A.)
@@ Acknowledgements.
The authors thank Samin Aref for useful comments and feedback. This work was supported in part by the US National Science Foundation under grants DMS–2005899 and DMS–2404617, and by computational resources provided by the Advanced Research Computing initiative at the University of Michigan.
@@ A.2 LFR network generation
The LFR networks we use for benchmarking are generated using the procedure described in [15], which is as follows.
@@
    Fix the number of nodes $n$ and mixing parameter $\mu$. In our studies we use node counts in the range $n\in[200,51200]$. The parameter $\mu$ controls the relative number of edges within and between communities. For small $\mu$ there are many more edges within communities than between them, which makes the communities relatively easy to detect. But as $\mu$ gets larger there are more edges between communities and detection becomes more difficult in a manner reminiscent of the detectability threshold in the standard stochastic block model [30, 34]. In our studies we use values of $\mu$ in the range $[0.2,0.8]$.
@@
    Draw a degree sequence from a power-law distribution with exponent $\tau_{1}$. Many networks have power-law degree distributions, typically with exponents between 2 and 3 [35], and the LFR model exclusively uses power-law distributions. We use $\tau_{1}=2.5$, with average degree $\langle k\rangle=20$ and maximum degree (which depends on graph size) $k_{\text{max}}=n/10$. Empirically, however, our results do not seem to be very sensitive to these choices.
@@
    Draw a set of community sizes from a power-law distribution with exponent $\tau_{2}$. Many networks are also found to have community sizes that approximately follow a power law, with typical exponents in the range from 1 to 2 [36, 37, 38, 15]. We use $\tau_{2}=1.5$ with a minimum community size of $s_{\text{min}}=20$ in all cases, while the maximum community size is set to $s_{\text{max}}=\max\left\{n/10,100\right\}$. Again, results were not particularly sensitive to these choices, provided they produce a valid distribution at all.
@@
The parameter values above are similar to those used for instance in [12]. As in that study, we find that algorithm performance is dictated primarily by the parameters $n$ and $\mu$, so it is these parameters that are varied our summary figures.
The LFR model is similar to a special case of the degree-corrected stochastic block model (DCSBM) [26], and hence one might expect that inference-based community detection methods employing the latter model would perform optimally on LFR networks. Specifically, in the limit of an infinite number of sampled networks and perfect optimization of each community detection method, the final performance measure for any algorithm is given by the expectation value of the similarity $M(g,h[A])$ between the ground truth LFR partition $g$ and the partition $h[A]$ of the LFR network $A$ inferred using the algorithm, where the expectation is taken over the ensemble $P(A,g|\theta)$ of LFR networks and partitions $A,g$ generated using parameters $\theta$ (meaning $\mu$, $\tau_{1}$, $\tau_{2}$, etc.). By using the LFR benchmark with parameters $\theta$ to compare the performance of community detection algorithms, we are therefore implicitly defining the “best” algorithm to be the one whose corresponding function $h[A]$ optimizes $\sum_{A,g}P(A,g|\theta)M(g,h[A])$. If the similarity measure we choose is the “all or nothing” error function $M(g,c)=\delta(g,c)$, then the optimal community detection algorithm is trivially the one with
@@
In other words, the optimal algorithm simply performs maximum a posteriori estimation under the model from which the network was generated. There is no explicit formula for the posterior probability under the LFR model, but to the extent that it is a special case of the DCSBM, we might expect the DCSBM (with appropriate priors) to give optimal results [39]. The LFR model, however, is not precisely a special case of the DCSBM. In particular, the DCSBM normally assumes a uniform distribution over community sizes, where the LFR model assumes a power law. Moreover, we are not using the crude all-or-nothing error function: our entire purpose in this paper is to develop mutual information measures that aggregrate and weigh different modes of error in a sensible fashion. These differences, it appears, are enough to ensure that the DCSBM does not perform the best in our testing.
Regardless, we emphasize that our use of the LFR benchmark in our analysis is simply for consistency with previous studies of network community detection methods [15, 10, 12]. The justification for our proposed similarity measure, on the other hand, is chiefly its theoretical merit over the conventional (symmetric, non-reduced) normalized mutual information, and is independent of the use of the LFR (or any other) benchmark.